### v1

# A Decade of News Forum Interactions: Threaded Conversations, Signed Votes, and Topical Tags

Abstract
We present a large-scale, longitudinal dataset capturing user activity on the online platform of DerStandard, a major Austrian newspaper. The dataset spans ten years (2013–2022) and includes over 75 million user comments, more than 400 million votes, and detailed metadata on articles and user interactions. It provides structured conversation threads, explicit up- and downvotes of user comments and editorial topic labels, enabling rich analyses of online discourse while preserving user privacy. To ensure this privacy, all persistent identifiers are anonymized using salted hash functions, and the raw comment texts are not publicly shared. Instead, we release pre-computed vector representations derived from a state-of-the-art embedding model. The dataset supports research on discussion dynamics, network structures, and semantic analyses in the mid-resourced language German, offering a reusable resource across computational social science and related fields.
Background and Summary
Social media have become a defining aspect of the first quarter of the $21^{st}$ century. For researchers, social media platforms are attractive data sources, as they offer access to large-scale, naturally occurring, and continuous streams of data on human communication and interaction. However, we increasingly find that straightforward access to such data is restricted, with many commercial platforms limiting or shutting down public APIs. Moreover, mainstream social platforms known for their discussions oriented at specific topics, such as political events or science, can quickly undergo rapid user migration or shifts in topical focus, as has been witnessed with Twitter/X [1]. In this context, online discussion forums attached to news media sites represent a valuable and relatively stable alternative. In prior work, news-based comment sections have been shown to influence readers’ perceptions of article quality [2, 3], to shape opinions [4], and to serve as counterpublic spaces that challenge dominant narratives [5]. These platforms typically combine structured comment threads, editorial context, and high-quality moderated textual content, making them particularly well suited for longitudinal and community-level analyses [6, 7].
We present a large-scale longitudinal dataset that provides user discussions on the online platform of DerStandard, an Austrian newspaper. Covering the full decade from 2013 to 2022, the dataset includes over 75 million user timestamped comments, more than 400 million up- and downvotes on these comments, and detailed metadata on the news articles, including editorial topic tags. A detailed overview of the file structure and metadata fields is provided in Figure 5.
DerStandard is an Austrian print newspaper that was founded in 1988. Already in 1995, it went online, being the first appearance of a major German-language newspaper on the web to their own claim. Its early adoption of digital community features, starting with chat rooms and evolving into discussion spaces for registered users below the news articles, has led to a highly active user forum. Users can post comments under articles, vote on others’ comments, and engage in structured discussion threads (see Figure 1 for an example of the interface). Moderation is performed using semi-automated systems that have been developed in collaboration with external research institutions such as the Austrian Research Institute for Artificial Intelligence (OFAI) [8], resulting in the minimization of low-quality text content on the platform. The user base of DerStandard is more male, younger and higher educated than the representative Austrian population [9, p. 3-4 in SI]
Figure 1: Example of the DerStandard comment platform interface. The screenshot shows a user comment with its associated voting summary in the upper right corner, indicating the number of upvotes (green) and downvotes (red). By hovering over the bar, users can see a breakdown of who voted and how (up- or downvote). Comments are displayed in threaded format with timestamps and reply links. We blurred user names in this example. User can also share and flag a comment as inappropriate (by clicking on the share and flag icons respectively)
To protect user privacy, we anonymize all persistent identifiers such as user and comment IDs using a salted cryptographic hash (see Section 1.2), and the raw text of comments is not publicly distributed by us. To respect these constraints while still enabling research that requires the use of text, we pre-computed vector representations for each of the more than 75 million texts in the dataset that we publicly share. We use a state-of-the-art, specialised open embedding model “KaLM-embedding-multilingual-mini-v1” [^1] https://huggingface.co/HIT-TMG/KaLM-embedding-multilingual-mini-v1 from the Hugging Face Model Hub.
With this dataset, we go beyond typical (US)-English-centric contributions by providing a large-scale textual data source in German, a mid-resourced language spoken by around 100 million people worldwide. Furthermore, the user base of DerStandard predominantly uses the Austrian variant of Standard German, which makes this dataset also an excellent source for studying the subtle variations from other variants of the pluricentric German language [10]. The period covered by our data includes numerous contentious political events in Austria and the world in general. Despite being a small European country, Austria is often considered a model for broader sociopolitical trends in Germany and other Western European countries. Its location in Central Europe and historical role as a crossing point between Eastern, Western, and Southern Europe give it shared characteristics with many of its neighboring countries.
The data combines several features in a unique way. First, it has a strong longitudinal scope, covering a full decade in the first quarter of the 21st century. Second, the platform’s active voting mechanism on comments provides explicit information on the sign of user interactions (up- or downvotes). Third, we retrieved detailed metadata on the comment as well as the article level from the platform, including content tags. These features make the dataset particularly well-suited for advancing research on online human behavior, especially through the analysis of network structures [11], context-dependent interactions [12], thread conversations [13], and temporal dynamics [14]. Particularly, explicit agreement and disagreement interactions between users are usually not available to researchers, and proxy measures, such as Wikipedia edits [15] or Reddit replies [16], have often been used instead to infer positivity or negativity of interactions. The dataset also includes metadata on deleted users, enabling studies of user churn and moderation effects. Importantly, this dataset provides unique opportunities to study large-scale online interactions and can be readily extended to include future years by public webscraping.
Data from DerStandard has previously been included in publicly released annotated datasets. While these datasets are very valuable because of their manual annotations and released text content, they do not offer the possibility to study the structural properties of user interactions or discussion threads. The One Million Post Corpus[17] features over one million user comments, with a subset manually annotated for content moderation tasks. GERMS-AT[18] contains comments annotated for varying degrees of sexism and misogyny, developed for the GermEval2024 shared task on sexism classification at the Conference on Natural Language Processing (KONVENS)  [^2] See https://ofai.github.io/GermEval2024-GerMS/. The availability of these datasets highlights the platform’s ongoing relevance for research. The other publicly released data sets complement our work by providing annotated subsets that support supervised learning and evaluation tasks. We describe in detail the relation between the datasets, including overlap and integration challenges, in Section 3.2.
The data shared in this article has already been used in three separate studies. The first work provided near real-time monitoring of the affective expressions of the entire DerStandard community during COVID-19 [19]. Second, it has been shown that the results of sentiment analysis on users’ comments on DerStandard closely track the dynamics of explicitly expressed mood by users in a traditional survey that was run on the platform in the same time period [9]. Third, information on vote interactions was used in order to extract signed networks of users and identify the main divisions on the platform. This analysis allowed to pinpoint issues that reinforce societal fault lines and contribute to polarization, as well as topics, such as COVID-19—that spark online conflict without strictly following these dividing lines [20]. In Section 3.2, we include the mapping of user IDs to the two factions identified in that study, enabling researchers to analyze behavior and interactions through the lens of long-term ideological alignment.
We provide technical validations for several aspects of our dataset. We confirm internal consistency of the provided data files and quantify potentially problematic cases such as deleted users or mismatched timestamps. In addition, we validate the embeddings that we share by comparing semantic similarity scores across reply structures and topic categories. We find that the embeddings encode both conversational structure and thematic coherence, making them suitable for semantic analysis for many research questions without having to access the full text.
Altogether, ”A Decade of DerStandard” provides a rare combination of scale, depth, and structure with the added benefit of a non-English setting. It is a valuable resource for research domains such as computational social science, network analysis, and natural language processing research.
1 Methods
1.1 Data Acquisition
We retrieved all information in the dataset by sending standard HTML requests to the public platform at derstandard.at. We used a combination of scripts, building on the curl library. We took care to balance the load on the platform with waiting times and added information to the user header that identified our requests as automated. All the information that we retrieved is publicly available to any visitor of the site, i.e. accessible without being logged into a user account on the platform. We downloaded raw HTML files and used XPath with xidel as well as jq for JSON processing to extract relevant objects. We also scraped article metadata, for example we extracted the article-associated topic tags annotated by DerStandard (see Figure 2). We converted all extracted data fields to a common TSV format.
Figure 2: Most discussed topics across DerStandard’s internal tag hierarchy. The diagram shows the internal tag hierarchy used by DerStandard editors to classify articles, highlighting the tags associated with the highest total number of user comments between 2013 and 2022. Flows between columns indicate how general categories are structured into more specific topics, proportional to the number of comments going to each subtopic. Node size indicates the volume of comments in each caWe color nodes according to the top-level category that they belong to. The ”other’ category represents any top- or mid-level tags that are not in the diagram due to having significantly lower comment counts than the ones displayed.
1.2 Anonymization
To protect user privacy while preserving relational links within the dataset, we anonymized all original identifiers of users and comments derived from the platform with new unique values created through a salted cryptographic hashing function. We used the BLAKE2s hashing algorithm [21] with an 8-byte digest and a fixed, undisclosed salt value. Original identifiers were converted to strings, concatenated with the salt, and hashed to produce consistent but irreversible anonymized values. This approach ensures that the same anonymized, hashed ID always maps to the same original identifier across all files, allowing for linkage (e.g., from comments to users or from votes to voters) without exposing sensitive information. Article identifiers were not anonymized, so they can be used to access the corresponding articles directly on the DerStandard platform (by appending the Article_Id at the end of https://derstandard.at/story/Article_Id).
1.3 Preparation of Additional Metadata: Thread Information and Users Dataset
To support research that requires computationally intensive operations over the full dataset, we pre-compute and provide several forms of aggregated metadata, for example, discussion thread characteristics. A thread refers to a structured sequence of comments starting from a top-level post and including all its nested replies. For each comment, we identify its thread origin parent, i.e. the top-level comment in the thread to which it belongs, as well as its thread depth, defined as the number of reply steps separating it from the root comment. These fields, together with the direct parent, allow a faster reconstruction of thread structures and the analysis of conversational hierarchies. For specific research questions, this provided information allows researchers to do an initial sampling of the data (for example, by considering threads of a certain depth only). We show two example thread discussions and their corresponding reply and vote networks in Figure 3.
We create a comprehensive users file containing one row per user who has either posted a comment or cast at least one vote. This file includes user-level aggregates such as the total number of comments written, total upvotes and downvotes received and given, and timestamps of first and last recorded activity (voting or commenting). These aggregated user profiles facilitate analyses of user behavior and engagement over time, with the opportunity of pre-selecting a subset of engaged and active users. To illustrate this, we showcase a signed network of a pre-selected subset of highly active users (high activity both in time and volume of votes) in Figure 4, which features a polarized and clustered structure that could not be observed by representing the whole set of users in the network.
Figure 3: Example forum conversations from three different perspectives. We illustrate two thread discussions on the platform using three visualizations: (I) the thread structure, showing the hierarchical relation between comments (with comments by the original comment author in green); (II) the reply network among users (highlighting the original post author in green); and (III) the vote network between users, including both commenting users (in yellow) and users who only voted (in gray), as well as the original author (in green). Example A shows a larger, mostly positive discussion, while Example B features a smaller, more decentralized discussion with a polarized voting pattern.
Figure 4: Network of active users. This network represents user-user voting interactions over the full decade, thresholded to include only users who gave or received at least 100 votes per year on average and were active for at least 3000 days. Edges represent positive (blue) and negative (red) voting interactions and are displayed only if their weight (number of positive or negative votes between a pair of users) exceeds 100 votes. Node size corresponds to the total user-weighted degree in the network (sum of displayed positive and negative votes). Node color encodes a positivity index, ranging from red (–1, fully negative) to blue (+1, fully positive), calculated as the difference between the number of positive and negative connections a user has, normalized by their total number of connections. A highly active cluster dominated by positive interactions emerges on the top-left part of the network. Two larger clusters, connected by more negative interactions between them but internally more positive, occupy the center and right side of the network. This figure is obtained by applying the GEPHI Force Atlas 2 layout algorithm to the positive network and adding the negative edges afterwards.
1.4 Text Embeddings
We generate embeddings of each comment in our dataset to include in the public release instead of the full text to protect the privacy of their content and avoid maintaining a public copy of, in the meanwhile, possibly deleted comments. While prior work has shown that embeddings can leak information, these methods rely on specific conditions that do not apply in our case. For example, some require access to both pretrained and fine-tuned embeddings to exploit their differences [22], or rely on generative models trained on auxiliary data to produce paraphrased versions of the input [23]. To date, no work has demonstrated that the exact original texts can be reliably recovered from plain fixed embeddings like those we release.
We provide dense vector representations with 896 dimensions for each comment using the “KaLM-Embedding” [24] approach. We selected this specialised class of models for the task because they generally perform well in benchmarks such as the Massive Text Embedding Benchmark [25]. Specifically, we used the ”KaLM-embedding-multilingual-mini-v1” [^3] https://huggingface.co/HIT-TMG/KaLM-embedding-multilingual-mini-v1 model from the Hugging Face Model Hub  [^4] Commit ID of the exact model version used: 8a82a0cd2b322b91723e252486f7cce6fd8ac9d3. The manageable size of around 2 GB of the model and smaller model complexity compared to many other approaches, allowed us to create a CPU-only workflow on the high-performance computing infrastructure available to us. We split the corpus of comments into arbitrary chunks (24 chunks for each of the 120 months of our data) to create a highly parallelizable setup for model inference.
2 Data Records
Figure 5: Overview of the data records included in the DerStandard dataset. The dataset is structured into several files based on record type and temporal granularity. A single Users file contains anonymized metadata per user, including activity statistics and voting behavior. Comments, Votes, and Embeddings are provided as monthly files, capturing all user-generated content, up- and downvotes, and pre-computed text embeddings, respectively. Articles files are aggregated yearly and include metadata such as title, URL, timestamps, and up to three editorial topic labels per article. All variable names coincide with the column header in the files. Additional explanations of some variables are provided in parentheses.
The full dataset is shared as a Dataset inside the Dataverse A Decade of DerStandard Forum Interactions [26]  [^5] https://dataverse.bsc.es/citation?persistentId=perma:BSC/P32CXW, organized into separate directories by data type and temporal resolution. These include user-level information, comment-level content, voting behavior, and article metadata, as well as pre-computed text embeddings. The files are stored in compressed TSV format for accessibility and ease of use. Except for the user file, all data is provided in either monthly or yearly resolution to facilitate efficient loading and period-specific analysis. An overview of the different data components and their variables is shown in Figure 5. The folders in the dataset, including data components and other summary and helper files, are the following:
  • Articles Files (Yearly)
  • Comments Files (Monthly)
  • Users Files (Single file)
  • Votes Files (Monthly)
  • Embeddings (Monthly)
  • Summary Files, addressing data inconsistencies, derived summary metadata and mapping to previous work annotations.
  • Scripts, containing basic example code for accessing and processing the data (e.g. to retrieve summary files)
Relationships between records across files can be established through shared identifiers. For instance, comment files link to the user file via the Author_Id and User_Id fields, and each comment is associated with the corresponding news article through the Article_Id. All users—whether they appear as authors of comments, voters or both—are included in the user file. In addition, every comment can be linked to its corresponding embedding vector using the Comment_Id. As detailed in Section 1.2, all user and comment identifications are anonymized for privacy protection. A separate file with a list of comments with small inconsistencies is also provided in the folder. Refer to Section 3.3 for a detailed explanation of these inconsistencies and the presence of empty fields (i.e. ”NA” values).
In total, the dataset covers over $75$ million comments, $400$ million votes, and more than $580$ thousand articles, spanning a ten-year period from 2013 to 2022. A brief summary of dataset size and user activity is presented in Table 1.
Table 1: Summary statistics of the DerStandard dataset. We provide key metrics on the number of articles, user comments, votes, and users between 2013 and 2022, including averages, standard deviations, and activity indicators.
Statistic Value
Articles Total 586 942
Per year (average) 58 694
Per year (stdDev) 10 726
Number of unique Topic_1 labels 25
Number of unique Topic_2 labels 305
Number of unique Topic_3 labels 1 722
Comments Total 75 644 850
Per month (average) 6 303 738
Per month (stdDev) 171 829
Number of threads 21 957 527
Number of replies ($\sim 71\%$) 53 617 134
Embeddings Vector dimension 896
Votes Total 412 511 165
Total positive ($\sim 77\%$) 319 154 131
Per month (average) 3 437 590
Per month (stdDev) 1 000 560
Users Total 247 863
% That voted more than 10 times $\sim$52%
% That commented more than 10 times $\sim$36%
% Active for at least 1 year $\sim$44%
% Active for 10 years $\sim$0.45%
3 Technical Validation
3.1 Quality Validations of Embeddings
To validate that the pre-computed text embeddings accurately capture the semantic structure of user comments, we conducted a series of validation analyses. These analyses confirm that the embeddings preserve meaningful relationships between comments at both conversational and topical levels.
Figure 6: Validation of comment embeddings based on discussion structure. Cosine similarity distributions for different types of comment pairs. Similarity is highest for direct reply pairs, followed by comments within the same thread, and then comments under the same article (excluding direct replies). Cross-thread and random pairs show the lowest similarity, indicating that semantic proximity in the embeddings aligns with structural proximity in discussions.
We first evaluate whether the cosine similarity between comment embeddings reflects structural proximity in discussion threads (Figure 6), decreasing from direct replies (most proximal) to pairs from different threads and articles (most distant). Despite considerable variability in similarity scores across comment pairs, the average similarities decrease consistently with structural proximity: highest for reply pairs, followed by non-reply comments within the same thread, comments under the same article, and lowest for cross-thread and random pairs, which serve as a baseline. These results indicate that the embeddings capture semantic properties of comments that are consistent with the structure of discussion threads.
Figure 7: Semantic coherence of topic-specific comment embeddings. Left: Average pairwise cosine similarity among 1M pairs of comment embeddings within and across selected editorial topics, compared against a random baseline. Higher within-topic similarity (diagonal) indicates that the embeddings capture topic-specific semantic content, particularly for focused topics such as football. Cosine similarity is computed after subtracting the mean embedding of $1.2$M random comments to remove shared semantic structure. Asterisks (*) denote significance based on a two-sample t-test against the random self-similarity distribution ($p<0.001$). Blue (red) in the matrix boxes indicates a positive (negative) similarity. Right: Two-dimensional t-SNE projection (after PCA reduction to 50 dimensions) of 5,660 comment embeddings across four selected topics. Clear clustering is observed for focused topics, and spatial proximity between topics reflects the similarity patterns observed in the left panel.
We next examine whether comments associated with the same editorial topic exhibit semantic cohesion (Figure 7). We compute cosine similarity after subtracting the mean embedding of $1.2$M random comments to remove shared semantic structure. We also compare all topic sets with a random set and use the self-similarity of the random set as a baseline distribution to compare significance (based on a two-sample t-test). Within-topic similarities (diagonal) are consistently higher than cross-topic similarities, with more focused topics such as refugees and football showing stronger internal coherence, while broader categories like policy issues display slightly lower cohesion. When comparing different topics, we observe that semantically related topics (e.g., refugees and policy issues or refugees and Middle East) exhibit higher cross-topic similarity, whereas unrelated topics (e.g., football and Middle East) show lower similarity levels, even below the comparison to the random set.
As a qualitative illustration of our embedding validation, the t-SNE projection on the right of Figure 7 provides a view of topic clustering in the embedding space. We first reduced the 896-dimensional comment embeddings to 50 principal components using Principal Component Analysis (PCA), retaining approximately 24% of the variance, and then applied t-SNE with a perplexity parameter of 50. The resulting visualization shows that comments associated with each topic form distinguishable clusters, with football forming the most compact and semantically distinct cluster.
Together, these analyses illustrate that the KaLM-based embeddings capture semantic relationships that reflect both conversational and topical structures, suggesting their potential usefulness for tasks such as topic classification, semantic retrieval, and comment clustering.
3.2 Mapping to Previous DerStandard Datasets
DerStandard has previously been used to create and share datasets supporting a variety of research goals. The One Million Post Corpus released by OFAI [17] contains over one million user comments, with a manually annotated subset targeting categories relevant to content moderation. By matching exact timestamps, we identified $1,415$ of these annotated comments in our dataset. Unmatched comments may fall outside our time frame or may have been removed by users or moderators, as the original annotations focused primarily on inappropriate or discriminatory content.
The GERMS-AT dataset [18], developed for the GermEval 2024 shared task on sexism classification at the Conference on Natural Language Processing (KONVENS) [^6] See https://ofai.github.io/GermEval2024-GerMS/, includes approximately $8,000$ DerStandard comments annotated for varying degrees of sexism and misogyny. Although based on the same platform, GERMS-AT does not include metadata or identifiers that would enable linkage with our dataset.
While integration with previously annotated corpora is limited, we provide a user-level mapping  [^7] Unique File Identifier: UNF:6:usMKWTPpkBEITjWROwLjVg== to the two main factions identified in prior work on affective polarization using DerStandard data [20]. This study applied the FAULTANA pipeline to detect disagreement fault lines on the platform by constructing a signed user-user network based on patterns of positive and negative interactions (i.e., comment votes). The method identifies an optimal partition into user factions by minimizing the number of frustrated edges—those that contradict the assumption of within-group agreement and across-group disagreement.
Using this approach, two main factions were detected, reflecting a persistent division in the user community over the 2014–2021 period. We include these faction labels as a user-level attribute for the $14,817$ users categorized in the study, with approximately 62% belonging to the larger faction. These labels provide a structural grounding for future research on ideological alignment, polarization dynamics, and cross-cutting issue detection. Note that the mapping is static and reflects long-term alignment; it does not account for potential shifts in user position over time due to future events or evolving platform dynamics.
3.3 Internal Consistency
Missing values occur in certain fields due to user deletions. We keep this data because it might be useful for research that studies user abandonment or migration in discussion-based platforms. On average, $8.46\%$ of comment authors are missing (Author_Id = NA) in the comments files, with a standard deviation of $3.78\%$ across months. Similarly, $8.33$(±$\pm 3.66$)of author fields are missing in the votes files, and $7.27$(±$\pm 3.35$)of voter identifiers are absent.
To ensure internal coherence of the dataset, we conducted a series of validation steps across all files and identifiers. We first verified that all Comment_Id values appearing in the votes files are also present in the comments files, ensuring referential integrity between user votes and the associated content.
We highlight a few anomalies and provide an explanation on how these issues have been dealt with in the published dataset: We keep the data and provide a separate file  [^8] Unique File Identifier: UNF:6:IWJ/35fdoMHY59b093o9mg== with a list of annotated Comment_Id with respective inconsistencies, followed by a column indicating the issue.
Approximately $1,000$ article identifiers referenced in the comments files do not have corresponding entries in the articles’ metadata. These cases are predominantly from older discussions, and the majority of them are inaccessible on the DerStandard website. We identify and list all $382,167$ comments ($0.5\%$ of the total amount of comments) belonging to these problematic article identifiers within the category of Article deleted.
In some cases, the number of votes recorded in the votes files for a given comment is one higher than the number of votes reported in the comments files. This discrepancy affects only a small subset of comments and is not systematically associated with deleted authors, thread structure, or specific articles. It appears to result from minor inconsistencies in how vote totals were recorded on the platform. To resolve this, we replaced the vote counts in the comments files with the corresponding aggregated values from the votes files.
Finally, we observed a small number of comments ($1,208$) with timestamps that precede the timestamp of the article they are linked to. This issue is also visible on the DerStandard website and may stem from article republication, delayed timestamping, or legacy content migration. One such example is available at https://www.derstandard.at/story/3222966/kosovo-unabhaengigkeit-und-separatistische-bewegungen. These rare cases may slightly affect the computation of lifetime estimates of a very small subset of users (i.e. less than $0.15\%$ of the $247,864$ users) or temporal dynamics and should be considered carefully. We list all comments with inconsistent time stamps within the category of Inconsistent timestamp.
Usage Notes
As noted in the main text, we provide a separate file listing all comment identifiers associated with potential data inconsistencies. These include: (1) comments referencing article identifiers that are no longer available in the articles metadata (nor accessible on the DerStandard website), and (2) comments with timestamps preceding those of their associated articles. While these cases represent a small fraction of the data, they may affect specific analyses, such as estimating user lifespans. We recommend that users assess whether these entries impact their workflow and, if necessary, exclude the small number of affected rows from their analysis.
When working with article text, we advise combining the title and subtitle fields into a single string. In some instances, one of the two fields may be missing or empty; merging them helps preserve as much article context as possible.
Detailed instructions for downloading, decompressing, and using the dataset are included in the main README file within the repository. We strongly recommend that users consult this file before accessing the data. For example, the comment embedding vectors are split into multiple parts to avoid exceeding file size limits ($2.5$ GB per file). These files must be concatenated before decompression.
Code Availability
Supporting code and example scripts to process the data can be found in the dataset folder [26]. Example code to join and decompress the embeddings files from the repository can be found in the README file.
We do not release the original web scraping code used to collect the data, as it involves platform-specific workarounds that may not apply at the current time and could place unnecessary strain on the servers of DerStandard. By sharing the processed dataset directly, we aim to reduce redundant scraping activity and minimize the load on the platform’s infrastructure.
Acknowledgements and Funding
We thank Elena Candellone, Darja Cvetković, Özgür Togay, David March, Timur Naushirvanov, Simon D. Lindner, and Elena G. de Lamo for testing the dataset during the Winter Workshop on Complex Systems 2025 and providing valuable feedback on its integrity and usability.
E.F. and V.G. acknowledge support from the project CNS2022-136178 financed by MCIN/AEI/10.13039/501100011033 and by the European Union Next Generation EU/PRTR.
This work is part of Maria de Maeztu Units of Excellence Programme CEX2021-001195-M, funded by MICIU/AEI /10.13039/501100011033.